Skip to main content

Change Notifications

View the Global HSC Change Calendar 

3/30 - Moodle Upgrade 6:00 pm - 8:00 pm

On Thursday, March 30, 2017, HSC IT will be performing an upgrade from version 3.1.2 to version 3.1.5 and as a result Moodle will be down from 6:00pm to 8:00pm MT.

4/9 - New HSC Firewall to be installed 10:00 am - 8:00 pm

HSC Network will be installing a new Firewall on Sunday, April 9, from 10am to 8pm. During this time there may be short interruptions in internet service. This will affect both outgoing internet traffic such as web browsing and incoming traffic such as VPN and Citrix from outside the HSC network.

Network Security will be monitoring for issues caused by the changeover. Any issues believed to be firewall related after the follow-up announcement should be entered into Help.HSC

TBD - Move Poison Center Voice Recording Server

Plan to move the voice recording server HigherGround and the VoIP lines in the Poison Center call centers to a new network stack.

3/19 - Cerner Outage 11:00 am - 3:00 pm

  • ALL CERNER APPLICATIONS WILL BE DOWN: PowerChart, SurgiNet, FirstNet, PM Office, Scheduling, Radnet, etc., will be down during the upgrade window.
  • NO SCANNED DOCUMENTS WILL BE AVAILABLE (this includes scanned EKG's, PFT's, scanned clinic logs, etc.)
  • ALL SYSTEM INTERFACES WILL BE DOWN.
  • The voice dictation applications will be available. Dictated reports will not be available in PowerChart until after the downtime. Radiology reports for orders placed during the downtime will be available only in the iSite PACS as an "annotation" on an image until after the conclusion of the downtime.
  • Critical labs will be called to floor. For lab results questions, contact Tricore Service Center @ 938-8922.
  • Electronic prescription orders, eligibility checks, and medication history requests from ePrescribing will be down.
  • Exchange of patient demographic and clinical information between Cerner Millennium and the following systems will be down: Sunquest lab orders & results; PowerPath AP results; ePrescribing prescriptions, eligibility & med history; eScription transcriptions; Radiology systems (iSite PACS, PowerScribe, PenRad mammography, Primordial); PyxisRx medication cabinets; CBORD dietary orders; Lumedx cardiac cath; Midas+ quality management; Pyxis Supply materials management; RxExpress outpatient pharmacy; SMS billing; IDX billing; 3M chart coding; CRTC Velos eResearch; New Mexico Health Information Exchange (NMHIC HIE) clinician's portal; SUnRx; and TeleTracking bed management.
  • Cerner downtime registration will be available on the UNMH IntraNet after the downtime starts.
  • Cerner’s downtime application (724Access) will be available after the downtime starts. Link is located on UNMH IntraNet web page under Web Based Systems, Cerner 724 Downtime Application. Tip sheet located on the SynergE3 Intranet website at: https://hospitals.health.unm.edu/intranet/synerge3/index.shtml.

3/19 - Network Outage HSLIC 8:00 am - 2:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections. Outages will be around 6 hours and will be done mostly on a weekend morning to minimize impact

3/21 - Drive Mapping Change

On Tuesday March 21, 2017 the HSC CIO’s office (Academic and Research) will be converting access to shared drives from script based mapping to group policy based mappings. Drive letters should remain unchanged. This change is the result of a Microsoft Windows Update that no longer allows symbolic links between mappings and has left several workstations without mapped drives after a user logs on to the workstation. If you have workstations that will not map drives after Wednesday March 22, and after a reboot, then please enter a service request in Help.HSC.

3/25 - Network Outage FPC, Surge, 1001 Med Arts 8:00 am - 2:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections. Outages will be around 6 hours and will be done mostly on a weekend morning to minimize impact

3/11 - SharePoint Maintenacne 10:00 am - 6:00 pm

SharePoint will be down for maintenance on 3/11/2017 from 10:00 am to 6:00 pm.

3/12 - CANCELED New HSC Firewall to be installed 10:00 am - 8:00 pm

HSC Network will be installing a new Firewall on Sunday, March 12, from 10am to 8pm. During this time there may be short interruptions in internet service. This will affect both outgoing internet traffic such as web browsing and incoming traffic such as VPN and Citrix from outside the HSC network.

Network Security will be monitoring for issues caused by the changeover. Any issues believed to be firewall related after the follow-up announcement should be entered into Help.HSC

3/13 - Main Campus Firewall Changes 6:00 am - 7:00 am

On Monday, March 13, IT Data Networking Group will upgrade the edge firewall to the latest code to prevent manageability issues. During this change, the entire campus will see an outage lasting 5 to 10 minutes as all of the routers re-establish with their peers. All services in the IT Data Center, VPN, and external services will be unavailable.

Additionally, IT Data Network Group will be upgrading the firewalls in the IT Data Center in building 153. Customers will see a brief outage of a few pings lost durinig the change over period. No security policy is changing.

2/12 Upgrade Domenici Hall Distribution Switches 8:00 am - 10:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections.

2/11 Upgrade UPC Distribution Switches 2:00 pm - 7:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections.

2/8 - 2/12 Move CTH network from OSIS to Core

To make our long-term architecture work we need to have this connected to the two core switches instead of OSIS. This will be done after hours and one link at a time and should have no detectable impact to users or system.

2/3 - 2/4 Network Outage Health Sciences and Services Bldg. 7:00 pm - 6:00 am

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections.

1/28 Network Outage Reginald Fitz Hall 8:00 am - 2:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections.

1/22 Upgrade ACC floors 2,3,4, & 5 to 10Gb switches 6:00 pm - 2:00 am

For ACC 2, 3, 4 and 5 - We will migrate the switches to 10G and configure the uplinks into port-channels on the access switch side and aggregated ethernet ports on the distribution layer side. No disruptive work will occur until the clinic closes.

1/22 Network Outage Family Practice Center 10:00 am - 6:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections.

1/17 Upgrade HSC Wireless Network Controller 6:00 pm - 7:00 pm

Upgrade HSC-WISM-S7 controller from IOS 8.0.133.0 to 8.2.130.0 in order to support the CISCO 2800 APs being installed. This will impact the north campus HSC education and research buildings only. There may be intermittent wireless outages during this time.

1/17 - 1/20 Various SharePoint Outages (See Details)

HSC IT will need to perform maintenance on the SharePoint farm. This change will address several issues with the system and it will result in the following outages:

Tuesday January 17, 2017 9:00 pm - 11:00 pm
Wednesday January 18, 2017 6:00 am - 8:00 am
Wednesday January 18, 2017 9:00 pm - 10:00 pm
Friday January 20, 2017 9:00 pm - 11:00 pm

1/14 Help.HSC System Upgrade 8:00 am - 5:00 pm

To solve several outstanding issues, and to improve performance. Help.HSC will be updated to version 8.3.1 on 1/14/2017. The service will be un-available from 8:00 AM - 5:00 pm

1/14 UNMH Netscaler Upgrade 9:00 am - 10:00 pm

UNMH IT will be upgrading Citrix Netscaler on Saturday January 14th to improve performance and stability. Citrix will also present an updated look to the web page, but applications in Citrix will work the same. There may be brief interruptions in Citrix apps during the upgrade. Will also apply to other Netscaler services like Outlook and the Hospital intranet, but users should not notice those changes. Each service will be moved individually and tested before moving to the next.

Just a note that a redirect has been created and tested so that anyone going to http://receiver.health.unm.edu or https://receiver.health.unm.edu will be redirected to https://citrixweb.health.unm.edu to be put in place at the time of the change.

1/7 Network Outage Biomedical Research Facility 8:00 am - 2:00 pm

HSC IT and UH Networking are in the process of replacing obsolete building network access switches, which connect all network devices (computers, printers, VOIP phones and etc.) to the building distribution switches. This process will cause interruption in service for wired and wireless network connections.

12/14 HSLIC/COP Student Printing Services Outage 8:00 am - 1:00 pm

Student printing services for HSLIC and the COP lounge will not be available Wednesday morning in order to move the pay per print system service to a new provider.

12/7 Install New Blades in UNMH Data Center 8:00 pm - 10:00 pm

Add new blades to the Nexus switches to support expansion. After the blades are installed, the uplinks for these switches will be moved form the old core switches to the new core switches

12/3 - 12/4 - Deploy new Internet Edge Firewall

This is a replacement for FWSM currently embedded in the HSC edge routers. Sporadic temporary interruptions to Internet service may occur.

11/30 - Rolling Outage of VOIP Phones 5:30 pm - 10:00 pm

UNM IT will be doing maintenance on the VOIP phone system. There will be rolling outages of VOIP phones in order for each of the to reboot.

11/30 Install New Blades in BBRP Switch 8:00 pm - 10:00 pm

Add new blades to the Nexus switches to support expansion. After the blades are installed, the uplinks for these switches will be moved form the old core switches to the new core switches

Both of these operations are in-service changes and should not result in any downtime.

11/6 - iSite Upgrade to Intellispace 6:00 pm - 6:00 am

PACS iSite will be upgraded to a new version this weekend. The data migration will actually start Sunday at 6:00 pm and be ready by Monday morning at 6:00 am. A new client is required to access Intellispace. The old client will be used up to Sunday at 6:00 pm. Both clients can be installed on a workstation at the same time

The installers are located at http://10.235.64.176/installers and the server address it will need is isitepacs.health.unm.edu. You may install on workstations now, but it will not work until Monday after 6:00 am. After it is live the installers will also be located at isitepacs.health.unm.edu.

11/6 - Replace Vassar Offices Network Distribution Layer 4:00 pm - 10:00 pm

Replace the 3550's that make up the distribution layer of all Vassar Offices with a pair of QFX-5100s.

Estimated actual outage should be fairly short, but user impact may be noted from fiber plant changes needed before the actual switch over.

11/2 - BBRP Data Center Work

This change will move the network racks about 8 feet to integrate with the rest racks in the BBRP data center. In addition, two new racks will be installed. No network cables will be disconnected during the move, and power will be moved one side at a time. In the event of catastrophic power loss, all equipment in the network racks will fail over to other data centers. The switching in the existing racks is not dependent on the racks being moved